# Relaxation of quasi-convex functionals with variable exponent growth

### v2

@@ 3. BV-type spaces
Let $p:\Omega\to[1,\infty)$ be lower semicontinuous. Then
@@
Assume first that $u\in{\rm BV}(\Omega)\cap W^{1,p(\cdot)}(\Omega\setminus Y)$ and $\varrho_{\text{\rm old}}(u)<\infty$. If $\varrho_{\varphi^{*}}(|w|)<\infty$, then $\varphi^{*}(x,|w|)=\infty\chi_{(1,\infty)}(|w|)=0$ for almost all $x\in Y$, so that $|w|\leqslant 1$ a.e. on $Y$. Thus
$\begin{split}\varrho_{V,{p(\cdot)}}(u)&=\sup\bigg\{\int_{\Omega}u\mathop{div}\nolimits w\,dx-\int_{\Omega\setminus Y}\frac{|w|^{p^{\prime}(x)}}{p^{\prime}(x)}\,dx\,\Big|\,w\in C^{1}_{0}(\Omega,{\mathbb{R}}^{n}),\varrho_{\varphi^{*}}(|w|)<\infty\bigg\}\\ &\leqslant\sup\bigg\{\int_{\Omega}u\mathop{div}\nolimits w\,dx-\int_{\Omega\setminus Y}\frac{|w|^{p^{\prime}(x)}}{p^{\prime}(x)}\,dx\,\Big|\,w\in C^{1}_{0}(\Omega,{\mathbb{R}}^{n}),|w|\leqslant 1\text{ a.e. on }Y\bigg\}\end{split}$
@@ 4. Proof of main result
Let $\Omega$ be a bounded open set with Lipschitz boundary and $p\in C(\overline{\Omega})$. If $f:\mathbb{R}^{m\times n}\to[0,\infty)$ satisfies (H0), (H1), (H2) and (H3), then the functional $F_{qc}:{\rm BV}^{p(\cdot)}(\Omega;{\mathbb{R}}^{m})\to[0,\infty)$, defined as
@@
Let $p$ be $\log$-Hölder continuous and $f:{\mathbb{R}}^{n\times m}\to[0,\infty)$ satisfy (H0), (H2) and (H3). Then

### v3

@@ 2. Preliminaries
We will need the following technical lemma. Note that we do not need the $\log$-Hölder continuity in all of $\Omega$, only the stronger version in the set $Y$, so we say that a variable exponent $p$ is strongly $\log$-Hölder continuous in $Y$ if
$p(x)-1\leqslant\frac{\omega(|x-y|)}{\log(e+\frac{1}{|x-y|})},$
Lemma 2.4.
Suppose that $p$ is strongly $\log$-Hölder continuous in $Y$. If $\varrho_{p^{\prime}(\cdot)}(w)<\infty$ for some continuous $w$, then $|w|\leqslant 1$ in $Y$
Proof.
Suppose that $|w(y_{0})|>1$ for some $y_{0}\in Y$. Let $B\subset\Omega$ be a ball centered at $y_{0}$ such that $|w|>c_{0}>1$ in $B$. By the strong $\log$-Hölder continuity condition in $Y$,
$p^{\prime}(x)\geqslant\frac{1}{p(x)-1}\geqslant M\log(e+\frac{1}{|x-y_{0}|}),$
where $M$ can be made large by restricting $x$ to a small ball centered at $y$. Consequently,
$\int_{\Omega}|w|^{p^{\prime}(x)}\,dx\geqslant\int_{B}c_{0}^{M\log(e+\frac{1}{|x-y_{0}|})}\,dx\geqslant\int_{B}|x-y_{0}|^{-M\log c_{0}}\,dx,$
where in view of [15, Lemma 3.1.6] the left hand side is equivalent to $\varrho_{p^{\prime}(\cdot)}(w)$. In a small ball centered at $y_{0}$, $M\log c_{0}\geqslant n$ and the integral diverges. Thus, if the integral converges, then $|w|\leqslant 1$, as claimed. ∎
@@ 3. BV-type spaces
Let $p:\Omega\to[1,\infty)$ be lower semicontinuous and strongly $\log$-Hölder continuous in $Y$. Then
@@
Assume first that $u\in{\rm BV}(\Omega)\cap W^{1,p(\cdot)}(\Omega\setminus Y)$ and $\varrho_{\text{\rm old}}(u)<\infty$. If $\varrho_{\varphi^{*}}(|w|)<\infty$, then $|w|\leqslant 1$ on $Y$ by Lemma 2.4. Thus
$\begin{split}\varrho_{V,{p(\cdot)}}(u)&=\sup\bigg\{\int_{\Omega}u\mathop{div}\nolimits w\,dx-\int_{\Omega\setminus Y}\frac{|w|^{p^{\prime}(x)}}{p^{\prime}(x)}\,dx\,\Big|\,w\in C^{1}_{0}(\Omega,{\mathbb{R}}^{n}),\varrho_{\varphi^{*}}(|w|)<\infty\bigg\}\\ &\leqslant\sup\bigg\{\int_{\Omega}u\mathop{div}\nolimits w\,dx-\int_{\Omega\setminus Y}\frac{|w|^{p^{\prime}(x)}}{p^{\prime}(x)}\,dx\,\Big|\,w\in C^{1}_{0}(\Omega,{\mathbb{R}}^{n}),|w|\leqslant 1\text{ on }Y\bigg\}\end{split}$
@@ 4. Proof of main result
Let $\Omega$ be a bounded open set with Lipschitz boundary and $p$ be strongly $\log$-Hölder continuous. If $f:\mathbb{R}^{m\times n}\to[0,\infty)$ satisfies (H0), (H1), (H2) and (H3), then the functional $F_{qc}:{\rm BV}^{p(\cdot)}(\Omega;{\mathbb{R}}^{m})\to[0,\infty)$, defined as
@@
Let $p$ be strongly $\log$-Hölder continuous and $f:{\mathbb{R}}^{n\times m}\to[0,\infty)$ satisfy (H0), (H2) and (H3). Then
@@
Acknowledgements The last author is a member of GNAMPA-INdAM whose support is gratefully acknowledged. She is also indebted with Gianni Dal Maso and Ana Margarida Ribeiro for helpful discussions on the subject of this paper.